# Leveraging Auxiliary Task Relevance for Enhanced Industrial Fault Diagnosis through Curriculum Meta-learning

### v1

@@ II-A2 Related Task-aware Meta-learning
Directly transferring patterns from the auxiliary tasks to the target task may introduce noise thus adversely affecting the model performance. By holding the principle of “paying more attention to more relevant knowledge”, we further consider the relevance of different working conditions in different tasks when conducting the global update. We obtain the tasks relevance (e.g., $\gamma_{rel}$) based on different vibration signal patterns, we thus attentively adapt the gradient across tasks by employing their relevance.
@@
where indices $i$ designate different auxiliary tasks, $t$ is the target task and $k$ traverses the latent dimensions. The derived distances, namely $\gamma_{rel_{0,t}}$, $\gamma_{rel_{1,t}}$, and $\gamma_{rel_{2,t}}$, serve as task relevance, with values closer to 1 suggesting more related tasks. In other words, if the auxiliary task is more related to the target task, we adapt the gradient so that it updates faster in that direction. There, the Eq.(1) is reformulated as:
@@ II-A3 Curriculum Strategy
In light of the challenges presented by high task diversity, we introduce a novel strategy inspired by the curriculum learning paradigm to enhance the convergence rate and generalization capacity of the meta-learner. We advocate for a sequential presentation of tasks to the meta-learner, adhering to an “easier first, harder later” doctrine. To operationalize this, we introduce a teacher recommender. For each auxiliary task, the recommender undergoes preliminary training and uses the highest validation score as an indicator of task difficulty. This metric subsequently informs the design of a curriculum for task sampling at each meta-training iteration. More specific, to quantify the difficulty of a given task, we partition the samples into a training set $D_{\text{train}}$ and a validation set $D_{\text{valid}}$. We train a base LSTM model extensively on $D_{\text{train}}$ as tearch recommender. Using the performance metric $\Phi$ (e.g., AUC) on the validation set, we ascertain the task-level difficulty, denoted as $\delta_{c}$. Specifically, $\delta_{c}$ is inversely related to the highest validation score achieved over all model parameters $\theta$:
@@
II-B Target-level Fine-tune
Specifically, assuming the model (e.g., LSTM) through meta-level training contains $L$ layers, we freeze first $l$ $(1\leq l\leq L)$ layers, while adding $n$ layers after the $l$ layers and fine tune the parameters by using data of the target task, so as to maximize the reusability of the general parameters. Hence, $\mathbf{e}_{hist}^{(tar)}$ is fed into the recurrent layer to infer the hidden state $\boldsymbol{h^{u}_{t_{k}}}$ at $t_{k}$, given by,
$\boldsymbol{h^{u}_{t_{k}}}=LSTM_{frozen}(\mathbf{e}_{hist}^{(tar)}).$ (7)
By doing this, the freezing-layer operation helps generate a network, capable of better balancing the parameters between the auxiliary tasks and target task after the final model fine-tuning.
$\boldsymbol{\hat{y}}=softmax(f(\mathbf{h}^{u}_{t_{k}}))$ (8)
where $\mathbf{y}$ is an one-hot embedding of the ground-truth fault. Algorithm 1 summarizes the training process of CMEFD, consisting of meta training (lines 3-11), freezing layers and model fine tuning (lines 12-14), as well as fault diagnosis (lines 15-16).
@@ II-B Target-level Fine-tune
9:        Sample another $N$ vibration signal from $\mathbb{D}_{i}$ as the eval_batch based on the difficulty;
@@ III Case Study
TABLE I: COMPARISON OF TRADITIONAL DL-BASED AND META-LEARNING-BASED MODELS
(*The best results are highlighted in bold.)
@@
Rolling bearings play a vital role as essential components in electric machines and find extensive usage across various mechanical equipment. Given the prolonged rotational wear and exposure to high temperatures during operation, bearing failures are common faults that occur. Therefore, conducting research on fault diagnosis in rolling bearings holds significant importance. This study specifically focuses on the detection of faults in rolling bearings within electric motor, considering the varying working conditions and a limited number of available samples. To evaluate the effectiveness of the proposed model, a series of case studies are conducted using two widely recognized datasets: the Case Western Reserve University (CWRU) Bearing Dataset [25], the Paderborn University Rolling Bearing Dataset [26] and the Gearbox dataset [19].
Moreover, several state-of-the-art models have been applied for comparison, including LSTM [23], the capsule network (Cap-Net) [4], wide deep CNN (WDCNN) [27], MAML [17], WDCNN-few-shot (WCFS) [10], CHAML [20], PKAML [19]. The LSTM, Cap-Net and WDCNN are DL-based model, MAML, WCFS, CHAML and PKAML are meta-learning based models.
(a) Healthy
(b) Ball 0.007
(c) Inner Race 0.007
(d) Outer Race 0.007
Figure 2: The visualisation of the vibration signals corresponding to different fault types under same fault diameter and working condition.
III-A Case 1: CWRU Bearing Dataset
III-A1 Dataset Description
The CWRU dataset has been commonly used to verify DL fault diagnosis methods. The selected rolling CWRU bearing dataset encompasses a comprehensive range of fault types, as depicted in Table II. The dataset comprises four distinct types of electric machine health states, which encompass one normal state and three fault states. The sizes of the faults are specified in terms of diameters, measuring 0.007, 0.014, and 0.021 inches. Fig. 2 provides a visual representation of the vibration signals corresponding to different fault types, all with the same fault diameter. As mentioned in Section II, fault diagnosis under various working conditions can be regarded as different tasks, and each task is a multi-class classification problem. Detailed descriptions of each task and relationship between the motor load and motor speed shown are presented in Table III. We treat one of the tasks as the target task and the rest as auxiliary tasks each time. Following [28], we chronologically divide the dataset of the target task into training, validation, and test sets with a ratio of 8:1:1.
@@ 1 Healthy 0
III-A2 Hyper-parameter Settings
The optimal hyper-parameter settings for all models are empirically found out based on the performance on the validation set. Specifically, for TL-based moded, the learning rate is selected from {0.1, 0.05, 0.01, 0.005, 0.001, 0.0001}, and batch size is set from {256, 128, 64}. For meta-learning based models, the learning rate $\alpha,\beta$ are searched from {0.1, 0.01, 0.001, 0.0001}; and the batch size is set from {256, 128, 64} for a fair comparison. For RT-ACM, the number of freezing laysers $l$ is searched in range of [1,4] stepped by one, where the best setting is 3 for all tasks.
Figure 3: Accuracy in the (a) outer loop and (b) inner loop.
Figure 4: Model performance on $T^{1}$.
III-A3 Performance Comparison
@@
Table I presents a comprehensive comparison of traditional deep learning-based models and meta-learning-based models across various tasks and shots. The results indicate a clear performance superiority of meta-learning-based models over their traditional counterparts. Specifically, our method RT-ACM consistently achieves the highest accuracy across all scenarios, notably outperforming other models with an impressive accuracy range of 98.51% to 99.73% in the 3-way tasks and 93.22% to 97.48% in the 6-way tasks. MAML, WCFS and CHAML, while also meta-learning-based, demonstrate strong performances, especially in the 3-way 5-shot settings. CHAML generally outperforms WCFS due to its structured learning process. By leveraging a curriculum sampling, CHAML progressively exposes the model to increasingly complex tasks, enhancing its ability to generalize and perform well across diverse scenarios. This staged learning process allows the model to build a more robust understanding of the task at hand, resulting in higher accuracy, particularly in more challenging 6-way tasks. Our proposed method RT-ACM integrates related task-aware meta-learning with a curriculum strategy, which further amplifies its performance. The related task-aware mechanism ensures that the model is adept at identifying and leveraging similarities between tasks, thereby improving its efficiency in learning new tasks. Additionally, the curriculum strategy within RT-ACM helps in systematically increasing the difficulty of the tasks presented during training, allowing the model to develop a deeper and more nuanced understanding of the task space. This dual approach not only enhances the learning process but also ensures that the model remains adaptable and highly accurate, as evidenced by its superior performance across all evaluated metrics.
@@
Fig. 3a shows the performance comparison of four meta-learning-based models, in terms of accuracy over increasing global update steps. Our proposed method, RT-ACM, which leverages related task-aware meta-learning and a curriculum strategy, demonstrates a superior ability to rapidly achieve high accuracy. This rapid improvement can be attributed to the efficient transfer of knowledge from related tasks and a structured learning process that progresses from simpler to more complex tasks. In contrast, CHAML, which employs curriculum sampling, lags behind RT-ACM and fails to match its performance throughout the training process due to a less effective related task prioritization. WCFS initially shows competitive accuracy but experiences a decline in performance in later stages, possibly due to overfitting or poor generalization. MAML, while maintaining steady progress, does not achieve the same level of accuracy as RT-ACM. Fig. 3b illustrates the relationship between the performance of model and the number of local update steps. Through empirical analysis, it is observed that a single-step update is adequate to enhance the accuracy significantly.
Fig. 4 illustrates that the RT-ACM model stands out for its exceptional combination of efficiency and accuracy on 3-way 5-shot. When compared to conventional models like LSTM and WDCNN, RT-ACM’s testing time is considerably shorter,which reflecting its higher processing speed. Additionally, RT-ACM competes closely in accuracy with more sophisticated models like MAML and WCFS. These results highlight the potential of RT-ACM, underscoring the advantages of incorporating auxiliary tasks to enhance the target task in terms of accuracy. Moreover, the curriculum transfer of knowledge from auxiliary tasks in RT-ACM demonstrates rapid learning capabilities.
(a) LSTM
(b) MAML
(c) WCFS
(d) RT-ACM
Figure 5: The visualization of classification from four models.
(a) LSTM
(b) MAML
(c) WCFS
(d) RT-ACM
Figure 6: The visualization of confusion matrices from four models.
III-B Case 2: PU Bearing Dataset
@@ III-B1 Dataset Description
Different from CWRU dataset, the Paderborn University (PU) Bearing Dataset is a pivotal diagnostic tool for rolling bearings include the artificially damaged data. To mitigate the influence of similar bearings on the dataset, vibration signals were collected from 26 bearings characterized by artificial damage, natural damage, and healthy states. The detailed working conditions are presented in Table IV. Following [19], the first 20 bearings were utilized for training, while the remaining 6 bearings were allocated for testing. Each bearing exhibited four distinct working conditions are presented in Table V, primarily varying in rotational speed, load torque, and radial force.
TABLE IV: INFORMATION OF USED BEARINGS FROM PU DATASET
No. Bearing Code Faulty Type Damage Method
@@ 4 K006 Healthy -
5 KA01 Outer Race Muliple Damage
@@ 9 KA08 Outer Race Drilling
10 KI01 Inner Race Muliple Damage
@@ 26 KI18 Inner Race Fatigue
TABLE V: DETAILED DESCRIPTIONS OF EACH TASK UNDER PU DATASET
@@
TABLE VI: THE COMPARATIVE METHODS IN PU DATASET ($T^{1}$)
(*The best results are highlighted in bold.)
Models Accuracy Precision F1-score
10-way
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
LSTM 40.15 50.89 39.92 50.05 39.61 49.84
MAML 76.85 85.35 76.71 85.19 76.63 85.12
WCFS 78.52 91.07 78.37 90.92 78.45 91.03
CHAML 78.75 90.20 77.61 91.07 78.58 91.13
PKAML 79.20 92.09 79.13 91.97 79.16 92.01
RT-ACM 81.57 94.34 81.43 94.21 81.46 94.27
(a) Training Loss
(b) Validation Accuracy
Figure 7: The visualization of training loss and validation accuracy from four models.
III-B2 Evaluation Metrics
Given the inclusion of artificially damaged data in the PU dataset, resulting in imbalanced distribution, precision and F1 score offer better insights into model performance. Thus, three established evaluation metrics are employed [6]: Accuracy, F1-Score and Precision. Accuracy quantifies the proportion of correct predictions in the total predictions made. It is given by the formula:
Accuracy=Number of Correct PredictionsTotal Number of $\text{Accuracy}=\frac{\text{Number of Correct Predictions}}{\text{Total Numberof Predictions}}$
Accuracy provides a general measure of model performance, especially when the class distributions are balanced. Precision is the fraction of relevant instances among the retrieved instances. In other words, it quantifies the accuracy of positive predictions. It is given by:
Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$
Precision provides insight into the reliability of positive predictions made by the model. The F1 score is the harmonic mean of precision and recall. It balances the trade-off between precision and recall and is especially useful when the class distribution is imbalanced. It is computed as:
[…]
Where, precision is the proportion of true positive predictions in the set of positive predictions made. Recall is the proportion of true positive predictions in the set of actual positives.
III-B3 Performance Comparison
@@
Table VI shows a performance comparison of various models on the PU dataset in 10-way, 1-shot and 5-shot. Notably, all methods except LSTM are meta-learning based approaches, allowing them to handle data sparsity issues more effectively than LSTM.. Among these methods, PKAML employs a pretraining process to integrate prior knowledge, enhancing its performance. CHAML utilizes a curriculum sampling in meta-training, providing a structured approach to learning. However, both PKAML and CHAML do not adequately consider the diversity and relevance of auxiliary tasks in relation to the target task, which can limit their effectiveness. In contrast, our proposed method, RT-ACM, incorporates related task-aware meta-learning along with a curriculum strategy, resulting in superior performance. This is evidenced by RT-ACM achieving the highest scores across all metrics and shots, with an accuracy of 81.57% and 94.34% in 1-shot and 5-shot settings, respectively, outperforming the runner-up PKAML and other models significantly.
Fig. 5 presents t-SNE visualizations for four models, highlighting their ability to discriminate between classes. The LSTM visualization shows poorly separated clusters, indicating significant class overlap and lower classification efficacy. MAML improves with tighter clusters but still suffers from overlap at the cluster centers. WCFS slightly betters class separation, with clearer clusters but some outliers. RT-ACM’s visualization stands out, with each class forming distinct, isolated clusters, showcasing superior class discrimination and robust classification. The color-coded clusters denote various classes, and RT-ACM’s close-knit clusters suggest strong internal similarity and external dissimilarity among classes, reflecting RT-ACM’s effective feature space transformation for enhanced class distinction.
@@
Fig. 6 showcases confusion matrices for four classification model, illustrating each model’s predictive performance. The LSTM matrix shows widespread misclassifications, such as confusing class 2 with class 3, indicating difficulty in distinguishing similar classes. MAML improves, concentrating predictions along the diagonal but still mixes up classes like 4 and 5. WCFS further enhances accuracy, with clearer diagonal values and fewer errors, notably between classes 7 and 8. RT-ACM stands out with a strong diagonal focus, demonstrating superior accuracy and minimal misclassification, especially distinguishing between closely related classes like 5 and 6, and 9 and 10, where other models struggle. This highlights RT-ACM’s robust predictive capability across diverse classes.
Fig. 7a illustrates the training loss of various models, including LSTM, MAML, WCFS, and RT-ACM. The LSTM exhibits significant fluctuations, indicating instability from a small dataset size and potential overfitting issues. In contrast, RT-ACM shows a quick and steady reduction in loss, outperforming MAML and WCFS in terms of learning efficiency and model regularization, suggesting a superior dataset fit. Fig. 7b compares validation accuracy, with RT-ACM achieving higher accuracy early and maintaining it, highlighting its robustness and generalization capability. LSTM’s accuracy fluctuates, while MAML and WCFS improve steadily but don’t match RT-ACM’s performance. RT-ACM excels in convergence speed and generalization, benefiting from task-aware meta-learning and a curriculum strategy.

### v2

@@ II-A2 Related Task-aware Meta-learning
Directly transferring patterns from auxiliary tasks to the target task can introduce noise, potentially degrading model performance. To address this, we adhere to the principle of “paying more attention on more relevant knowledge” and incorporate the relevance of working conditions across tasks during the global update. By calculating task relevance (e.g., $\gamma_{rel}$) based on variations in vibration signal patterns, we adapt gradients across tasks with attention to their relevance.
@@
where indices $i$ designate different auxiliary tasks, $t$ is the target task and $k$ traverses the latent dimensions. The derived distances, namely $\gamma_{rel_{0,t}}$, $\gamma_{rel_{1,t}}$, and $\gamma_{rel_{2,t}}$, serve as task relevance, with values closer to 1 suggesting more related tasks. In essence, if an auxiliary task exhibits greater relevance to the target task, the gradient is adjusted to facilitate faster updates in that direction. Consequently, Eq. (1) is revised as follows:
@@ II-A3 Curriculum Strategy
To address the challenges posed by high task diversity, we propose a novel approach inspired by the curriculum learning paradigm to improve both the convergence speed and the generalization ability of the meta-learner. This approach adopts a sequential task presentation strategy, following the principle of “easier tasks first, harder tasks later.” To implement this, we introduce a teacher recommender, which undergoes initial training for each auxiliary task and utilizes the highest validation score to estimate task difficulty. This difficulty metric guides the curriculum design for task sampling during each meta-training iteration. Specifically, to measure task difficulty, we split the samples into a training set $D_{\text{train}}$ and a validation set $D_{\text{valid}}$. A base LSTM model is trained extensively on $D_{\text{train}}$ to serve as the teacher recommender. The difficulty of a task is quantified using a performance metric $\Phi$ (e.g., AUC) evaluated on $D_{\text{valid}}$. The task difficulty, denoted as $\delta_{c}$, is inversely proportional to the highest validation score obtained across all model parameters $\theta$:
@@
9:        Sample another $N$ signal from $\mathbb{D}_{i}$ as the eval_batch based on the difficulty;
@@
II-B Target-level Fine-tune
After extracting general patterns from auxiliary tasks via the meta-learning paradigm, fine-tuning the model for the target task becomes essential [28]. By investigating the freezing of layers and then fine-tuning, the network generalizes better than one trained directly on the target dataset. This approach not only integrates knowledge from auxiliary tasks but also facilitates better adaptation to the target task [29].
Assuming the meta-trained model (e.g., LSTM) consists of $L$ layers, we freeze the first $l$ layers $(1\leq l\leq L)$, add $n$ new layers after the frozen ones, and fine-tune the parameters using target task data to maximize the reuse of general parameters. Consequently, the target task input $\mathbf{e}_{hist}^{(tar)}$ is passed through the recurrent layer to compute the hidden state $\boldsymbol{h^{u}_{t_{k}}}$ at time $t_{k}$, as expressed by:
$\boldsymbol{h^{u}_{t_{k}}}=LSTM_{frozen}(\mathbf{e}_{hist}^{(tar)}).$ (7)
By doing this, the freezing-layer operation facilitates the creation of a network that achieves a more effective balance between the parameters learned from auxiliary tasks and those optimized for the target task, following the final fine-tuning of the model.
$\boldsymbol{\hat{y}}=softmax(f(\mathbf{h}^{u}_{t_{k}}))$ (8)
where $\mathbf{y}$ is an one-hot embedding of the ground-truth fault. Algorithm 1 summarizes the training process of RT-ACM, consisting of meta training (lines 4-12), freezing layers and model fine tuning (lines 13-14), as well as fault diagnosis (lines 15-16).
@@ III Case Study
TABLE I: COMPARISON OF DL-BASED AND META-LEARNING-BASED MODELS
(*The best results are highlighted in bold; the runner up is underlined.)
@@
Rolling bearings are critical components in electric machines and are widely utilized in various mechanical systems. Due to prolonged rotational stress and exposure to high temperatures, bearing failures frequently occur during operation. As a result, research on fault diagnosis for rolling bearings is of paramount importance. This study concentrates on detecting faults in rolling bearings within electric motors, taking into account diverse working conditions and the limited availability of samples. To assess the performance of the proposed model, a series of case studies are carried out using two well-known datasets: the Case Western Reserve University (CWRU) Bearing Dataset [30] and the Paderborn University Rolling Bearing Dataset [31].
Additionally, several state-of-the-art models are employed for comparison, including the capsule network (Cap-Net) [6], wide deep CNN (WDCNN) [32], MAML, TMSL, CHAML, and PKAML. Among these, Cap-Net and WDCNN are deep learning-based models, while MAML, CHAML, and PKAML belong to the category of meta-learning-based approaches.
@@ 1 Healthy 0
III-A Case A: CWRU Bearing Dataset
III-A1 Dataset Description
The CWRU dataset is widely utilized to evaluate fault diagnosis methods [2, 15, 33]. The selected rolling bearing dataset from CWRU includes a comprehensive set of fault types, as shown in Table II. It covers four distinct electric machine health states, including one normal state and three fault states. The fault sizes are specified by their diameters, which measure 0.007, 0.014, and 0.021 inches. As outlined in Section II, fault diagnosis under different working conditions can be considered as separate tasks, with each task representing a multi-class classification problem. Table III provides detailed descriptions of each task, along with the relationship between motor load and motor speed. For each iteration, one of the tasks is treated as the target task, while the remaining tasks are used as auxiliary tasks. Following [34], we chronologically divide the dataset of the target task into training, validation, and test sets with a ratio of 8:1:1.
III-A2 Performance Comparison
@@
Table I presents a comprehensive comparison of deep learning-based models and meta-learning-based models across various tasks and shots. The results indicate a clear performance superiority of meta-learning-based models over their traditional counterparts. Specifically, our method RT-ACM consistently achieves the highest accuracy across all scenarios, notably outperforming other models with an impressive accuracy range of 98.51% to 99.73% in the 3-way tasks and 93.22% to 97.48% in the 6-way tasks. MAML and TMSL, while also meta-learning-based, demonstrate strong performances, especially in the 3-way 5-shot settings. Fig. 3(a) presents the 3-way 5-shot performance results on Task 1, while Fig. 3(b) illustrates the 6-way 5-shot performance on the same target task. It is evident that TMSL generally outperforms MAML, attributed to its advantage in sequence learning processes. Our proposed method RT-ACM integrates related task-aware meta-learning with a curriculum strategy, which further amplifies its performance. The related task-aware mechanism ensures that the model is adept at identifying and leveraging similarities between tasks, thereby improving its efficiency in learning new tasks. Additionally, the curriculum strategy within RT-ACM helps in systematically increasing the difficulty of the tasks presented during training, allowing the model to develop a deeper and more nuanced understanding of the task space. This dual approach not only enhances the learning process but also ensures that the model remains adaptable and highly accurate, as evidenced by its superior performance across all evaluated metrics.
@@
Fig. 4(a) shows the performance comparison of four meta-learning-based models, in terms of accuracy over increasing global update steps. Our proposed method, RT-ACM, which leverages related task-aware meta-learning and a curriculum strategy, demonstrates a superior ability to rapidly achieve high accuracy. This rapid improvement can be attributed to the efficient transfer of knowledge from related tasks and a structured learning process that progresses from simpler to more complex tasks. In contrast, TMSL, which employs k-nn based task evaluation, lags behind RT-ACM and fails to match its performance throughout the training process due to a less effective related task prioritization. MAML, while maintaining steady progress, does not achieve the same level of accuracy as RT-ACM. WDCNN initially shows competitive accuracy but experiences a decline in performance in next stages, possibly due to overfitting or poor generalization.
Fig. 4(b) demonstrates the superior performance of the RT-ACM model in both running time in each iteration and accuracy on the 3-way 5-shot task. While RT-ACM requires slightly more computational time compared to deep learning models like Cap-net and WDCNN due to its meta-learning process of acquiring learning capabilities through different tasks, it significantly reduces the time consumption compared to other meta-learning approaches like MAML and TMSL. This improvement in efficiency can be attributed to the curriculum strategy’s progressive learning approach, moving from simple to complex tasks. Furthermore, RT-ACM achieves competitive accuracy levels, highlighting its potential in balancing the trade-off between computational efficiency and model performance. The incorporation of auxiliary tasks has proven beneficial in improving accuracy on the target task.
Figure 3: Comparison of (a) 3-Way and (b) 6-Way 5-Shot Performance on $T_{1}$.
Figure 4: Accuracy in the (a) outer loop and (b) model performance on $T^{1}$.
III-B Case B: PU Bearing Dataset
@@ III-B1 Dataset Description
Different from CWRU dataset, the Paderborn University (PU) Bearing Dataset is a pivotal diagnostic tool for rolling bearings include the artificially damaged data. To mitigate the influence of similar bearings on the dataset, vibration signals were collected from 26 bearings characterized by artificial damage, natural damage, and healthy states. The detailed working conditions are presented in Table V. Following [8], 32 unique bearing fault types are evaluated across four distinct operating conditions. To reflect the complexity of real-world industrial environments and to enhance the precision of diagnostic outcomes, each fault type is considered a separate category under each working condition, resulting in a total of 128 fine-grained fault categories. From these, 100 categories are randomly chosen for training purposes, while the remaining 28 are reserved for testing. Further specifics of the working conditions are provided in Table VI.
TABLE IV: THE COMPARATIVE METHODS IN PU DATASET ($T^{1}$)
(*The best results are highlighted in bold; the runner up is underlined.)
Models Accuracy Precision F1-score
5-way
5-shot 6-shot 5-shot 6shot 5-shot 6-shot
Cap-Net 46.61 48.85 46.52 48.03 46.36 48.47
WDCNN 72.48 74.06 71.46 74.57 68.47 75.08
MAML 89.65 90.9 87.90 88.92 89.06 90.77
[…]
RTACM 94.34 95.65 94.21 95.42 94.27 96.18
TABLE V: Information of Bearings Used from PU Dataset
No. Bearing Code Fault Type Damage Method
@@ 4 K006 Healthy -
5 KA01 Outer Race Multiple Damage
@@ 9 KA08 Outer Race Drilling
10 KA09 Outer Race Drilling
11 KI01 Inner Race Multiple Damage
12 KI03 Inner Race Electric Engraver
@@ 14 KI07 Inner Race Electric Engraver
15 KI08 Inner Race Electric Engraver
16 KA04 Outer Race Fatigue
@@ 32 KI18 Inner Race Fatigue
TABLE VI: DETAILED DESCRIPTIONS OF EACH TASK UNDER PU DATASET
@@
III-B2 Performance Comparison
@@
Table IV shows a performance comparison of various models on the PU dataset in 5-way, 5-shot and 6-shot. Notably, all methods except Cap-Net and WDCNN are meta-learning based approaches, allowing them to handle data sparsity issues more effectively than traditional DL-based model. Among these methods, MAML employs a meta-training process to integrate different tasks, enhancing its performance. TMSL utilizes a k-nn based sequence sampling in meta-training, providing a structured approach to learning. However, both MAML and TMSL do not adequately consider the diversity and relevance of auxiliary tasks in relation to the target task, which can limit their effectiveness. In contrast, our proposed method, RT-ACM, incorporates related task-aware meta-learning along with a curriculum strategy, resulting in superior performance. This is evidenced by RT-ACM achieving the highest scores across all metrics and shots, with an accuracy of 94.34% and 95.65% in 5-shot and 6-shot settings, respectively, outperforming the runner-up TMSL and other models significantly. Fig. 5 presents t-SNE visualizations for four models, highlighting their ability to discriminate between classes. The Cap-Net visualization shows poorly separated clusters, indicating significant class overlap and lower classification efficacy. MAML improves with tighter clusters but still suffers from overlap at the cluster centers. TMSL slightly betters class separation, with clearer clusters but some outliers. RT-ACM’s visualization stands out, with each class forming distinct, isolated clusters, showcasing superior class discrimination and robust classification. The color-coded clusters denote various classes, and RT-ACM’s close-knit clusters suggest strong internal similarity and external dissimilarity among classes, reflecting RT-ACM’s effective feature space transformation for enhanced class distinction.
@@
Fig. 6 showcases confusion matrices for four classification model, illustrating each model’s predictive performance. The Cap-Net matrix shows widespread misclassifications, such as confusing class 3 with class 4, indicating difficulty in distinguishing similar classes. MAML improves, concentrating predictions along the diagonal but still mixes up classes like 4 and 9. TMSL further enhances accuracy, with clearer diagonal values and fewer errors, notably between classes 9 and 10. RT-ACM stands out with a strong diagonal focus, demonstrating superior accuracy and minimal misclassification, especially distinguishing between closely related classes like 4 and 10, where other models struggle. This highlights RT-ACM’s robust predictive capability across diverse classes.
Figure 5: The visualization of classification from four models on (10-way 5-shot).
Figure 6: The visualization of confusion matrices from four models (10-way 5-shot).
Figure 7: (a) The performance of models across varying numbers of local update steps and (b) the impact of the number of freezing layers.
Fig. 7(a) illustrates a sensitivity analysis of a meta-learning-based model across varying numbers of local update steps. The results indicate that a single-step update is sufficient to significantly enhance accuracy, demonstrating that MAML-based models can achieve notable improvements with minimal updates. In Fig. 7(b), the impact of the number of frozen layers on model performance is examined. The performance initially improves as the number of frozen layers increases, reaching a peak when three layers are frozen, after which performance slightly declines. This suggests that three frozen layers offer the optimal balance for performance. Too few frozen layers may lead to a loss of auxiliary information, while too many frozen layers limit the model’s ability to adapt to the target task, resulting in overfitting.